1. The manager of a store that specializes in selling tea decides to experiment with a new blend. She will mix some Earl Grey tea that sells for $5 per pound with some Orange Pekoe tea that sells for $3 per pound to get 100 pounds of the new blend. The price of the new blend is to be $4.50 per pound, and there is to be no difference in revenue from selling the new blend versus selling the other types. How many pounds of the Earl Grey tea and the Orange Pekoe tea are required?
I'm in pre-calculus and this was a problem that was given to us. I understand that it's a system of linear equations, but I just can't seem to set the 2 equations up right. I was thinking the two equations would be: x+y=100; 5x+3y=4.50 but it just doesn't come out right. Any help on how to set up the equations would be great! Thanks.

Answer by Mathtut(3670)

your very close except on the 2nd equation the new mixture needs to be set to
4.50(100) so x+y=100
...........and 5x+3y=4.50(100)
x then =75
y=25